# JARVIS-VLA: Post-Training Large-Scale Vision Language Models to Play Visual Games with Keyboards and Mouse

### v1

@@ Abstract
Recently, action-based decision-making in open-world environments has gained significant attention. Visual Language Action (VLA) models, pretrained on large-scale web datasets, have shown promise in decision-making tasks. However, previous work has primarily focused on action post-training, often neglecting enhancements to the foundational model itself. In response, we introduce a novel approach, Act from Visual Language Post-Training, which refines Visual Language Models (VLMs) through visual and linguistic guidance in a self-supervised manner. This enhancement improves the models’ capabilities in world knowledge, visual recognition, and spatial grounding in open-world environments. Following the above post-training paradigms, we obtain the first VLA models in Minecraft that can follow human instructions on over 1k different atomic tasks, including crafting, smelting, cooking, mining, and killing. Our experiments demonstrate that post-training on non-trajectory tasks leads to a significant 40% improvement over the best agent baseline on a diverse set of atomic tasks. Furthermore, we demonstrate that our approach surpasses traditional imitation learning-based policies in Minecraft, achieving state-of-the-art performance. We have open-sourced the code, models, and datasets to foster further research. The project page can be found in https://craftjarvis.github.io/JarvisVLA.
@@ 1 Introduction
Pretraining foundation models on large-scale, noisy internet datasets has become a mainstream approach in NLP and vision [18, 1, 43, 38]. The success of models like GPT and LLAMA [40, 35] has shown that large, capable language models can infer and execute tasks described by language prompts. However, this paradigm has yet to achieve similar success in the decision-making domain [50, 14]. In particular, while OpenAI’s Video Pre-Training (VPT) model [4] has attempted to apply a similar approach in Minecraft, it still relies heavily on imitation learning (IL) after collecting large-scale YouTube videos of human play. VPT’s approach of pretraining with imitation learning, followed by downstream supervised fine-tuning and reinforcement learning, made significant strides—culminating in the successful crafting of diamond tools, a key challenge in Minecraft [^1] Diamond tools are considered a grand challenge, with experienced human players taking up to 20 minutes (24,000 actions) to craft them..
@@
Despite this success, the reliance on next-action prediction in imitation learning limits the development of robust, multi-task decision-making abilities [6, 49, 39, 34]. Moreover, this pretraining paradigm struggles to generalize to unseen environments or tasks due to the intricacies of the interactions between observations and behavior, whereas language tokens are more standardized. To overcome these challenges, a new approach has emerged that leverages pretrained Vision Language Models (VLMs) for decision-making. These models, known as Visual Language Action models (VLAs), integrate language understanding with action generation and can be further enhanced through post-training on visual-language tasks [25, 54, 7]. A more detailed discussion can be found in Figure 1 (left) and Section 2.2.
@@
However, much like traditional imitation learning, current VLA approaches predominantly focus on action post-training. In these models, the learning objective is to generate correct actions based on large-scale cross-task imitation data. We propose that, in addition to action generation, understanding the environment and incorporating task-related knowledge could be equally important for achieving more flexible and generalizable decision-making. To this end, we introduce a novel training paradigm—Visual Language Post-Training (ActVLP)—which integrates visual-language tasks into the post-training phase of VLA models. Following the above paradigms, we obtain the first VLA models in Minecraft that can follow human instructions on over 1k different atomic tasks, including crafting, smelting, cooking, mining, and killing.
@@
Our contributions are as follows: (1) We pioneer the use of VLA in the open-world environment of Minecraft by introducing JARVIS-VLA, a powerful model achieving state-of-the-art performance in action-based decision-making. (2) We introduce the concept of Visual Language Post-Training and identify key visual-language guidance strategies that enhance decision-making. (3) We investigate the scaling laws of VLA models, demonstrating that expanding the scale of non-trajectory vision-language tasks during post-training leads to significant improvements in downstream task performance. (4) We open-source the code, models, and datasets to support further research in this area.
@@
Figure 2: Previous VLA methods usually directly use imitation learning to finetune original vision-language models on large-scale multi-domain decision-making datasets to predict the actions [25, 7]. Our ActVLP training pipeline includes three stages: 1) post-training language models on text-only world knowledge with next-token prediction supervised fine-tuning, 2) post-training both vision encoder and language models on multimodal vision-language alignment and spatial grounding datasets with next-token prediction supervised fine-tuning, and 3) post-training only language models on multi-modal instruction following datasets with imitation learning.
@@ 2.1 Model Structure
As illustrated in Figure 1, JARVIS-VLA employs an architecture similar to Llava [27] but with slight modifications. The structural framework, consists of several key components:: 1) Visual Encoder: A Vision Transformer [18] that processes raw image pixels and converts them into a sequence of fixed-size image patches. 2) Image Projection Module: A lightweight two-layer MLP that projects image patch embeddings into the same representational space as word embeddings. 3) Language Model Transformers [40, 3]: A powerful autoregressive language model that serves as the core of the system, facilitating multimodal reasoning and decision-making.
Unlike OpenVLA [25], our framework is designed for partially observable environments. To accommodate this, we adopt a non-Markovian architecture by incorporating a history of observation images within the prompt. This approach ensures that the model retains temporal context, which is crucial for tasks requiring multi-step reasoning and long-horizon decision-making. In our experiments, we employ Llava-Next [27] and Qwen2-VL [43] as base vision language models, as both models provide robust support for multi-image reasoning, enabling enhanced perception and contextual understanding.
Another key distinguishing feature of JARVIS-VLA compared to prior VLA models is the integration of an action decoder. This module is responsible for generating both discrete and continuous actions. For discrete actions, we consolidate related action dimensions into unified categories to reduce redundancy and improve efficiency. For continuous actions, we discretize the action space into bins, which are then mapped to discrete tokens. These tokens are subsequently appended to the vocabulary of the original foundation model, allowing the model to generate both textual and action-based outputs in a unified manner.
Instead of retraining the base VLM’s tokenizer, we adopt a strategy inspired by RT-2 [7], repurposing the least frequently used tokens from the language tokenizer’s vocabulary to represent action semantics. Specifically, we replace the 51 least-used tokens, allocating 22 tokens for mouse control (e.g., cursor movements) and 29 tokens for special keyboard inputs (e.g., function keys and command shortcuts), which can be found in Appendix A. We introduce no other modifications to the original VLM architecture to maintain model generalizability and ensure broad compatibility across different foundation models. This design choice allows JARVIS-VLA to be easily integrated with various pre-trained multimodal models while preserving their inherent capabilities.
@@ 2.2 Training Pipeline
However, learning world knowledge from action-labeled trajectory data is inherently challenging [4]. Moreover, the lack of large-scale action-labeled datasets makes it challenging to pretrain expansive models using only trajectory data [34].
@@
Stage III: Imitation Learning on Trajectories
In the final stage, we fine-tune the VLMs on trajectory data, requiring the model to mimic expert actions given textual instructions $x_{\text{ins}}$ and the current observation image $o_{t}\in\mathbb{R}^{H\times W\times 3}$. The imitation learning objective is defined as:
@@
where $\pi$ represents the learned policy, and $a_{t:t+\tau}$ denotes the future action chunk predicted from the current step. During this phase, vision-related modules remain frozen, while the language tokenizer is modified to incorporate action tokens, and the language transformer undergoes full-parameter fine-tuning. Additionally, we adopt action chunking to further enhance training efficiency and optimize trajectory learning [15].
@@ 2.3 Datasets
To support the ActVLP training pipeline, we constructed a large-scale multimodal dataset. This dataset includes both non-trajectory task datasets for post-training and trajectory datasets for downstream imitation learning. The non-trajectory datasets are divided into three categories: knowledge-based question answering, visual language alignment, and spatial grounding. These categories are designed to enhance the model’s decision-making capabilities before trajectory fine-tuning. For trajectory datasets, we collected over 7.4 million frames of Minecraft gameplay data, including expert actions from diverse sources such as human-playing [4], youtube videos, and existing agents [45].
@@
VLA training is constructed on a dataset of human gameplay trajectories, particularly from the OpenAI contractor dataset in Minecraft [4], which includes diverse tasks. We also incorporated an additional 3M rollout frames from VPT [4] and JARVIS-1 [45] agents. For structured GUI-based tasks like crafting and smelting, we synthesized 6.4M expert data entries to improve imitation learning. Representative examples of our datasets are shown in Figure 3, with further details in Appendix D.
@@ 3 Experiments
Q2: Is visual language post-training the true cause of the performance improvement?
@@ 3.1 Experimental Setup
We use Minecraft 1.16.5 as our experimental platform [23]. As an open-world game with a substantial knowledge base on platforms such as Reddit and wiki [21], Minecraft poses significant challenges to agents while simultaneously offering rich resources for research. To ensure fair comparisons, we align the action and visual observation spaces with those of human players [4]. Additionally, we hide information unavailable to human players as well, such as agent location and inventory stats.
Benchmark and Evaluation Metrics
We conduct evaluations using two broad benchmarks: (i) the agent’s capacity to interact with the Minecraft environment to complete tasks; and (ii) vision-language tasks (e.g., question answering, spatial grounding) designed to assess the VLM’s understanding of Minecraft-specific knowledge. For the instruction-following tasks, we adopt the MCU Benchmark [31], focusing on four categories—Mine Blocks, Kill Entities, Craft Items, and Smelt Items—that represent a wide range of typical game-play behaviors in Minecraft. Notably, Craft and Smelt require 2D GUI manipulation through the mouse (covering thousands of item categories), whereas Mine and Kill involve recognizing, navigating, and interacting with targets in a 3D environment. Each category contains at least 5 distinct tasks. For instance, the Mine Blocks category includes mining iron ore [Uncaptioned image] with a stone pickaxe, oak logs [Uncaptioned image] with bare hands, grass [Uncaptioned image], dirt [Uncaptioned image], and obsidian [Uncaptioned image] with a diamond pickaxe. Our evaluation set includes both simpler tasks (e.g., mining oak logs) and more complex ones (e.g., mining obsidian for over 10 seconds) that have proven challenging for prior state-of-the-art agents [12, 30]. We perform each task at least 30 times and report the success rate per task, as well as the average success rate within each category. To ensure fairness, maximum execution steps for selected tasks match those reported by Lin et al. [31]. For vision-language assessments, the task formulations are illustrated in Figure 3. We provide human-written ground-truth answers and employ an LLM-as-judge to evaluate the performance of various VLMs (GPT-4o, Llava, Qwen-VL, and our post-trained VLMs). Detailed information on these vision-language benchmarks and results can be found in Appendix E.
Training and VLA Configurations
. Our training pipeline follows the process described in Section 2.2: we first obtain a visual-language post-training intermediate model, then further train it on trajectory tasks to produce the JARVIS-VLA. We conduct experiments using two popular frameworks: Qwen2-VL [43] and Llava [27]. We develop a discretized action tokenizer specific to Minecraft, comprising 51 tokens that represent camera movements and button actions. We utilize the trl SFT Trainer [41] for finetuning and deploy the VLA with vLLM [26]. Training is carried out on 32 A800-80G GPUs, while inference runs on a single NVIDIA RTX 3090. Further training details are provided in Appendix B.
Baselines
We compare our model with: 1) VPT [4], including both the behavior cloning (VPT-BC) and reinforcement learning (VPT-RL) variants; 2) STEVE-1[30], a text-conditioned policy that combines VPT and MineCLIP[21] for instruction following; 3) GROOT [12], which uses video prompts as task instructions; and 4) MineDreamer [58], which leverages a vision-language model and a diffusion model to guide the STEVE-1 policy. Each method follows the default configuration provided in the MCU benchmark for a fair comparison.
@@
VPT-BC [4] 248M 0.15 0.38 0.33 0.55 0.35 0.44 0.30 0.50 0.45 0.41 0.10 0.00 0.05
VPT-RL [4] 248M 0.05 0.35 0.25 0.35 0.25 0.28 0.50 0.30 0.62 0.55 0.05 0.35 0.20
STEVE-1 [30] 248M 0.20 0.35 0.54 0.30 0.75 0.38 0.45 0.20 0.70 0.57 0.25 0.40 0.33
GROOT [12] 248M 0.56 0.40 0.67 0.50 0.50 0.52 0.45 0.35 0.25 0.40 0.35 0.25 0.30
MineDreamer [58] 7B 0.25 0.40 0.55 0.30 0.70 0.39 0.50 0.25 0.30 0.42 0.30 0.30 0.30
Qwen2-VL (orig) 7B 0.77 0.60 0.79 0.93 0.80 0.84 0.83 0.53 0.40 0.60 0.03 0.10 0.07
@@
Table 1: Evaluation results of different policies on Minecraft tasks. Each group includes multiple tasks (at least 5), and the Avg. column reports the average success rate within each group. Qwen2-VL, Qwen2-VL (IL) and JARVIS-VLA-Qwen2-VL represent the training on the original qwen checkpoint, post-training on only large-scale imitation learning trajectories, and post-trained on VLP intermediate model. Qwen2-VL (ActVLP) achieves the highest success rates across all task groups. We test JARVIS-VLA-Qwen2-VL on 1k different instructions and the rollout videos can be found in the project page.

### v2

@@ Abstract
Recently, action-based decision-making in open-world environments has gained significant attention. Visual Language Action (VLA) models, pretrained on large-scale web datasets, have shown promise in decision-making tasks. However, previous work has primarily focused on action post-training, often neglecting enhancements to the foundation model itself. In response, we introduce Act from Visual Language Post-Training (ActVLP), a novel training paradigm. ActVLP distinctively enhances the foundation model prior to action-specific tuning by first post-training it on a curated set of environment-specific visual and linguistic tasks using self-supervised learning. This initial stage significantly improves the model’s capabilities in world knowledge, visual recognition, and spatial grounding. Subsequently, this strengthened VLM undergoes action post-training via imitation learning on trajectory datasets. Following this paradigm, we develop JARVIS-VLA, the first VLA model in Minecraft that can follow human instructions on over 1k different atomic tasks, including crafting, smelting, cooking, mining, and killing. Our experiments demonstrate that our ActVLP paradigm leads to a significant 40% improvement over the best agent baseline on a diverse set of atomic tasks. Furthermore, JARVIS-VLA surpasses traditional imitation learning-based policies in Minecraft, achieving state-of-the-art performance. We have open-sourced the code, models, and datasets to foster further research. The project page can be found at https://craftjarvis.github.io/JarvisVLA.
@@ 1 Introduction
Pretraining foundation models on large-scale, noisy internet datasets has become a mainstream approach in NLP and vision (Dosovitskiy, 2020; Achiam et al., 2023; Wang et al., 2024b; Team et al., 2023). The success of models like GPT and LLAMA (Touvron et al., 2023; OpenAI, 2023) has shown that large, capable language models can infer and execute tasks described by language prompts. However, this paradigm has yet to achieve similar success in the decision-making domain (Yang et al., 2023; Cheng et al., 2024). In particular, while OpenAI’s Video Pre-Training (VPT) model (Baker et al., 2022) has attempted to apply a similar approach in Minecraft, it still relies heavily on imitation learning (IL) after collecting large-scale YouTube videos of human play. VPT’s approach of pretraining with imitation learning, followed by downstream supervised fine-tuning and reinforcement learning, made significant strides—culminating in the successful ObtainDiamond, a key challenge in Minecraft [^1] Diamond tools are considered a grand challenge, with experienced human players taking up to 20 minutes (24,000 actions) to craft them..
@@
Despite this success, the reliance on next-action prediction in imitation learning limits the development of robust, multi-task decision-making abilities (Brohan et al., 2022; Wu et al., 2023; Team et al., 2024; O’Neill et al., 2023). Moreover, this pretraining paradigm struggles to generalize to unseen environments or tasks due to the intricacies of the interactions between observations and behavior, whereas language tokens are more standardized. To overcome these challenges, a new approach has emerged that leverages pretrained Vision Language Models (VLMs) for decision-making. These models, known as Vision Language Action models (VLAs), integrate language understanding with action generation and can be further enhanced through post-training on visual-language tasks (Kim et al., 2024; Zhen et al., 2024). A more detailed discussion can be found in Figure 1 (left) and Section 2.2.
@@
However, much like traditional imitation learning, current VLA approaches predominantly focus on action post-training. In these models, the learning objective is to generate correct actions based on large-scale cross-task imitation data. We propose that, in addition to action generation, understanding the environment and incorporating task-related knowledge could be equally important for achieving more flexible and generalizable decision-making. To this end, we introduce a novel training paradigm—Vision Language Post-Training (ActVLP)—which integrates visual-language tasks into the post-training phase of VLA models. Following the above paradigms, we obtain the first VLA models in Minecraft that can follow human instructions on over 1k different atomic tasks, including crafting, smelting, cooking, mining, and killing.
@@
Our contributions are as follows: (1) We pioneer the use of VLA in the open-world environment of Minecraft by introducing JARVIS-VLA, a powerful model achieving state-of-the-art performance in action-based decision-making. (2) We introduce the concept of Vision Language Post-Training and identify key visual-language guidance strategies that enhance decision-making. (3) We investigate the scaling laws of VLA models, demonstrating that expanding the scale of non-trajectory vision-language tasks during post-training leads to significant improvements in downstream task performance. (4) We open-source the code, models, and datasets to support further research in this area.
@@
Figure 2: Previous VLA methods usually directly use imitation learning to finetune original vision-language models on large-scale multi-domain decision-making datasets to predict the actions (Kim et al., 2024; Brohan et al., 2023). Our ActVLP training pipeline includes three stages: 1) post-training language models on text-only world knowledge with next-token prediction supervised fine-tuning, 2) post-training both vision encoder and language models on multimodal vision-language alignment and spatial grounding datasets with next-token prediction supervised fine-tuning, and 3) post-training only language models on multi-modal instruction following datasets with imitation learning.
@@ 2.1 Model Structure
As illustrated in Figure 1, JARVIS-VLA employs an architecture similar to Llava (Li et al., 2024a) but with slight modifications. The structural framework, consists of several key components:: 1) Visual Encoder: A Vision Transformer (Dosovitskiy, 2020) that processes raw image pixels and converts them into a sequence of fixed-size image patches. 2) Image Projection Module: A lightweight two-layer MLP that projects image patch embeddings into the same representational space as word embeddings. 3) Language Model Transformers (Touvron et al., 2023; Bai et al., 2023): A powerful autoregressive language model that serves as the core of the system, facilitating multimodal reasoning and decision-making.
Unlike OpenVLA (Kim et al., 2024), our framework is designed for partially observable environments. To accommodate this, we adopt a non-Markovian architecture by incorporating a history of observation images within the prompt. This approach ensures that the model retains temporal context, which is crucial for tasks requiring multi-step reasoning and long-horizon decision-making. In our experiments, we employ Llava-Next (Li et al., 2024a) and Qwen2-VL (Wang et al., 2024b) as base vision language models, as both models provide robust support for multi-image understanding, enabling enhanced perception and reasoning.
Another key distinguishing feature of JARVIS-VLA compared to prior VLA models is the integration of an action decoder. This module is responsible for decoding both discrete and continuous actions. For discrete actions, we consolidate related action dimensions into unified categories to reduce redundancy and improve efficiency. For continuous actions, we discretize the action space into bins, which are then mapped to discrete tokens. These tokens are subsequently appended to the vocabulary of the original foundation model, allowing the model to generate both textual and action-based outputs in a unified manner.
Instead of retraining the base VLM’s tokenizer, we adopt a strategy inspired by RT-2 (Brohan et al., 2023), repurposing the least frequently used tokens from the language tokenizer’s vocabulary to represent action semantics. Specifically, we replace the 51 least-used tokens, allocating 22 tokens for mouse control (e.g., cursor movements) and 29 tokens for special keyboard inputs (e.g., function keys and command shortcuts), which can be found in Appendix A. We introduce no other modifications to the original VLM architecture to maintain model generalizability and ensure broad compatibility across different foundation models. This design choice allows JARVIS-VLA to be easily integrated with various pre-trained multimodal models while preserving their inherent capabilities.
@@ 2.2 Training Pipeline
However, learning world knowledge from action-labeled trajectory data is inherently challenging (Baker et al., 2022). Moreover, the lack of large-scale action-labeled datasets makes it challenging to pretrain expansive models using only trajectory data (O’Neill et al., 2023).
@@
Stage III: Action Post-Training for Interaction
The final stage of our pipeline is action post-training, where the VLM transitions from a passive observer to an active agent. During this action post-training, the VLM is fine-tuned on trajectory data ($\mathcal{D}$). The model learns to map textual instructions ($x_{\text{ins}}$) and current visual observations ($o_{t}\in\mathbb{R}^{H\times W\times 3}$) to action chunks $a_{t:t+\tau}$. The imitation learning objective is to maximize the likelihood of these expert actions:
@@
where $\pi_{\theta}$ is the learned policy parameterized by $\theta$. We employ action chunking as this technique promotes temporally coherent and improves training efficiency (Kim et al., 2025). In this stage, the vision-related modules remain frozen. while the language tokenizer is modified to incorporate action tokens, and the language transformer undergoes full parameter fine-tuning.
@@ 2.3 Datasets
To support the ActVLP training pipeline, we constructed a large-scale multimodal dataset. This dataset includes both non-trajectory task datasets for post-training and trajectory datasets for downstream imitation learning. The non-trajectory datasets are divided into three categories: knowledge-based question answering, vision language alignment, and spatial grounding. These categories are designed to enhance the model’s decision-making capabilities before trajectory fine-tuning. For trajectory datasets, we collected over 7.4 million frames of Minecraft gameplay data, including expert actions from diverse sources such as human-playing (Baker et al., 2022), youtube videos, and existing agents (Wang et al., 2024c).
@@
VLA training is constructed on a dataset of human gameplay trajectories, particularly from the OpenAI contractor dataset in Minecraft (Baker et al., 2022), which includes diverse tasks. We also incorporated an additional 3M rollout frames from VPT (Baker et al., 2022) and JARVIS-1 (Wang et al., 2024c) agents. For structured GUI-based tasks like crafting and smelting, we synthesized 6.4M expert data entries to improve imitation learning. This trajectory data was partitioned into two distinct subsets. First, over 100 random trajectories for each MCU benchmark task were allocated for targeted fine-tuning. The remaining data, approximately 10 billion tokens, was then utilized for action post-training. Representative examples of our datasets are shown in Figure 3, with further details in Appendix D.
@@ 3 Experiments
Q2: Is vision language post-training the true cause of the performance improvement?
@@ 3.1 Experimental Setup
We use Minecraft 1.16.5 as our experimental platform (Guss et al., 2019). As an open-world game with a substantial knowledge base on platforms such as Reddit and wiki (Fan et al., 2022), Minecraft poses significant challenges to agents while simultaneously offering rich resources for research. To ensure fair comparisons, we align the action and visual observation spaces with those of human players (Baker et al., 2022). Additionally, we hide information unavailable to human players as well, such as agent location and inventory stats.
@@
VPT-BC (Baker et al., 2022) 248M 0.15 0.38 0.33 0.55 0.35 0.44 0.30 0.50 0.45 0.41 0.10 0.00 0.05
VPT-RL (Baker et al., 2022) 248M 0.05 0.35 0.25 0.35 0.25 0.28 0.50 0.30 0.62 0.55 0.05 0.35 0.20
STEVE-1 (Lifshitz et al., 2024) 248M 0.20 0.35 0.54 0.30 0.75 0.38 0.45 0.20 0.70 0.57 0.25 0.40 0.33
GROOT (Cai et al., 2024c) 248M 0.56 0.40 0.67 0.50 0.50 0.52 0.45 0.35 0.25 0.40 0.35 0.25 0.30
MineDreamer (Zhou et al., 2024) 7B 0.25 0.40 0.55 0.30 0.70 0.39 0.50 0.25 0.30 0.42 0.30 0.30 0.30
Qwen2-VL (raw) 7B 0.77 0.60 0.79 0.93 0.80 0.84 0.83 0.53 0.40 0.60 0.03 0.10 0.07
@@
Table 1: Evaluation results of different policies on Minecraft tasks, Each group includes multiple tasks (at least 5), and the Avg. column reports the average success rate within each group. Qwen2-VL, Qwen2-VL (IL) and JARVIS-VLA-Qwen2-VL represent the training on the original qwen checkpoint, post-training on only large-scale imitation learning trajectories, and post-trained on VLP intermediate model. Qwen2-VL (ActVLP) achieves the highest success rates across all task groups.
Benchmark and Evaluation Metrics
We conduct evaluations using two broad benchmarks: (i) the agent’s capacity to interact with the Minecraft environment to complete tasks; and (ii) vision-language tasks (e.g., question answering, spatial grounding) designed to assess the VLM’s understanding of Minecraft-specific knowledge. For the instruction-following tasks, we adopt the MCU Benchmark (Lin et al., 2023), focusing on four categories—Mine Blocks, Kill Entities, Craft Items, and Smelt Items—that represent a wide range of typical game-play behaviors in Minecraft. Notably, Craft and Smelt require 2D GUI manipulation through the mouse (covering thousands of item categories), whereas Mine and Kill involve recognizing, navigating, and interacting with targets in a 3D environment. Each category contains at least 5 distinct tasks. For instance, the Mine Blocks category includes mining iron ore [Uncaptioned image] with a stone pickaxe, oak logs [Uncaptioned image] with bare hands, grass [Uncaptioned image], dirt [Uncaptioned image], and obsidian [Uncaptioned image] with a diamond pickaxe. Our evaluation set includes both simpler tasks (e.g., mining oak logs) and more complex ones (e.g., mining obsidian for over 10 seconds) that have proven challenging for prior state-of-the-art agents (Cai et al., 2024c; Lifshitz et al., 2024). We perform each task at least 30 times and report the success rate per task, as well as the average success rate within each category. To ensure fairness, maximum execution steps for selected tasks match those reported by Lin et al. (2023). For vision-language assessments, the task formulations are illustrated in Figure 3. We provide human-written ground-truth answers and employ an LLM-as-judge to evaluate the performance of various VLMs (GPT-4o, Llava, Qwen-VL, and our post-trained VLMs). Detailed information on these vision-language benchmarks and results can be found in Appendix E.
Training and VLA Configurations
. Our training pipeline follows the process described in Section 2.2: we first obtain a visual-language post-training intermediate model, then further train it on trajectory tasks to produce the JARVIS-VLA. We conduct experiments using two popular frameworks: Qwen2-VL (Wang et al., 2024b) and Llava (Li et al., 2024a). We develop a discretized action tokenizer specific to Minecraft, comprising 51 tokens that represent camera movements and button actions. We utilize the trl SFT Trainer (von Werra et al., 2020) for finetuning and deploy the VLA with vLLM (Kwon et al., 2023). Training is carried out on 32 A800-80G GPUs, while inference runs on a single NVIDIA RTX 3090. Further training details are provided in Appendix B.
Baselines
We compare our model with: 1) VPT (Baker et al., 2022), including both the behavior cloning (VPT-BC) and reinforcement learning (VPT-RL) variants; 2) STEVE-1(Lifshitz et al., 2024), a text-conditioned policy that combines VPT and MineCLIP(Fan et al., 2022) for instruction following; 3) GROOT (Cai et al., 2024c), which uses video prompts as task instructions; and 4) MineDreamer (Zhou et al., 2024), which leverages a vision-language model and a diffusion model to guide the STEVE-1 policy. Each method follows the default configuration provided in the MCU benchmark for a fair comparison.